OpenStudy (anonymous):

(x+1)/(2x+2) simplified =?

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

its 1/2 so don't even bother lol

OpenStudy (mertsj):

\[\frac{x+1}{2x+2}=\frac{x+1}{2(x+1)}=\frac{1}{2}\]